The ability to taste the bitter compound phenylthiocarbamide (PTC) is an autosomal dominant trait. The inability to taste PTC is a recessive condition. In a sample of 500 people, 391 have the ability to taste PTC and 109 do not. Calculate the frequency of:

A. the recessive allele. Express your answer using two decimal places.

B.the dominant allele. Express your answer using two decimal places.

C.the homozygous dominant genotype. Express your answer using two decimal places.

D.the homozygous recessive genotype. Express your answer using two decimal places.

E. the heterozygous genotype. Express your answer using two decimal places.
Answer #1

Gene frequency = number of population with the trait / total number of population

Let p be the dominant allele

Let q be the recessive allele

Hardy Weinberg equations:

1 = p2 + 2pq + q2

1 = p + q

Frequency of recessive condition (qq) = 109/ 500 = 0.218

q = 0.46

p = 1 - q = 1-0.46 = 0. 54

A. Frequency of recessive allele (q) = 0.46

B. Frequency of dominant allele (p) = 0.54

C. Frequency of Homozygous dominant genotype ( p2 ) =0.29

D. Frequency of Homozygous recessive genotype ( q2 )= 0.21

E. Frequency of Heterozygous genotype (2pq) = 2 * 0.54 * 0.46 = 0.50